Find the work done to empty the cylendrical tub of radius r and height h filled with fluid of density rho.


1
Expert's answer
2020-10-02T07:09:33-0400


Let us consider a cylindrical tub of radius "r" and height "h" filled with a fluid of density "\rho" .

Let us consider a small element of the cylinder of thickness "dx" at a depth "x" from the top of the tub.

The volume of this element is

"d V=Area\times height\\
d V=\pi r^2\times dx"

The mass of this element is "dm=volume \times density = \pi r^2 \rho\ dx"

The weight of this element is "\pi r^2 \rho g dx"

The work done to lift this amount of fluid to a height "x" and thus emptying the cylinder is

"dW=weight\times distance"

"dW=\pi r^2\rho g\ xdx"

To empty the whole cylinder in such way the total work done is

"W=\int _{0}^{h} \pi r^2\rho g \ xdx\\
W=\pi r^2\rho g \int _{0}^{h} xdx\\
W=\pi r^2\rho g \left[x^2/2\right]_0^h\\
W=\frac{1}{2}\pi r^2\rho g h^2"


Answer: The work done to empty the cylindrical tub is "\frac{1}{2}\pi r^2 \rho g h^2" .